OpenStudy (anonymous):

HELP!! ASAP!!! 5x + 5 - 8x -4 = -29

OpenStudy (solomonzelman):

\[5x + 5 - 8x -4 = -29\]\[ - 3x + 1 = -29\]\[-3x+1\color{red} {-1}=-29\color{red} {-1}\]\[-3x=-30\]\[-3x\color{red} {\div(-3)}=-30\color{red} {\div(-3)}\]\[x=10\]

OpenStudy (anonymous):

x=10